OpenStudy (anonymous):

true or false. the square root of x^2 = x is an identity

OpenStudy (anonymous):

false

OpenStudy (anonymous):

why

OpenStudy (anonymous):

Because you the equation is not satisfied for every number you plug into it.

OpenStudy (anonymous):

\[2^2\neq2\]

OpenStudy (anonymous):

ohhh sorry. I did not see the square root. Wait

OpenStudy (anonymous):

\[\sqrt{x^2}=x\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

is not and identity either.

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

\[\sqrt{(-1)^2}\neq-1\]

OpenStudy (anonymous):

Because the equation is not always true.

OpenStudy (anonymous):

This is an identity:\[\sin^2\theta+\cos^2\theta=1\]

OpenStudy (anonymous):

No matter what value of theta you choose. the equation is always true.

OpenStudy (anonymous):

ok thanks. how about this question: if f(x)=g(x) is an identity with domain of validity D, which of the following must be true? a) for any x in D, fx is defined b) for any x in D, gx is defined. c) for any x in D, fx=gx (you can choose all of them, two of them, one of them, none, etc whichever u think are true)

OpenStudy (anonymous):

all of them are true.

OpenStudy (anonymous):

in the case of your first question the domain of validity D = positive numbers and zero.

OpenStudy (anonymous):

and then it is a identity.

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

You're welcome. By the way which course is this question from?

OpenStudy (anonymous):

ohh ok. Have fun with that!